# Solving the Hubbard model with Neural Quantum States

### v1

@@ Abstract
Figure 5: Energy difference between vertical stripe and horizontal stripe. Kinetic energy, potential energy, and total energy results are all shown. We systematically study systems $L_{x}\times 8$ with $L_{x}$ ranging from 8 to 32, with $t^{\prime}=-0.2$ and $\delta=1/8$. We can find horizontal stripe has lower energy for all the systems (notice that for $8\times 8$ system, vertical and horizontal stripe are the same). Moreover, the results show the gain of energy in horizontal stripe is mainly from the kinetic energy.
Previous numerical studies on stripe order in the Hubbard and related models have been conducted on cylindrical geometries with DMRG  (?, ?, ?, ?, ?). However, such boundary conditions and their related finite-size effect can inherently favor the formation of vertical stripe orientations. To alleviate the possible boundary effect, we study systems with PBC in both directions. We systematically study systems with size $L_{x}\times 8$ where $L_{x}$ is ranged from $8$ to $32$, with $t^{\prime}=-0.2$ and $\delta=1/8$ (see Equation 1). We consistently observe that a horizontal stripe configuration is energetically favored in the ground state while the system can accommodate half-filled stripe in both horizontal and vertical directions (see Figures 5S15-S20 in the Supplementary Materials, and Figure 4 (B) and (C)). These results suggest the existence of a tendency toward fewer long stripes rather than multiple fragmented ones in the systems, different from previous studies on cylinders  (?). From Figure 5, we can also find the gain of energy for the horizontal stripe state is mainly from the kinetic energy, while the potential energy in the horizontal stripe is slightly higher. A simple explanation is that the formation of long “river” of holes in the horizontal stripe is beneficial to the kinetic energy. A comprehensive understanding of this observation needs further investigations.
For rectangular systems with a width not divisible by $8$, a horizontal stripe with wavelength $\lambda=4$ is impossible, so “frustration” appears. An such example is the $16\times 10$ system, for which we find that the trend for horizontal stripe wins and the ground state has a wavelength of $\lambda=5$ while the vertical stripe state with $\lambda=4$ has a higher energy (see Figure S23-S24 in the Supplementary Materials). In another example for the $16\times 12$ system, the trend for wavelength $4$ wins in the competition and the ground state is a vertical stripe (see Figure S25 in the Supplementary Materials).
For the pure Hubbard model, it is also possible that the true ground state is horizontal stripe. But the resolution of this issue requires study of systems which can accommodate stripe with wavelength $8$ in both vertical and horizontal directions. The smallest size for such system under PBC is $32\times 16$ which is beyond our present capacity.
@@
In this work, we propose a new NQS framework for studying quantum many-body systems by leveraging the cutting-edge machine learning architectures and developing highly accurate optimization algorithms. State-of-the-art results for the 2D doped Hubbard model with unprecedented sizes are achieved with this new framework. NQS enables the study of 2D Hubbard model under periodic boundary conditions which is crucial to get rid of the boundary effect and makes the extrapolation to thermodynamic limit reliably. By studying 2D Hubbard model with large systems under periodic boundary conditions, we confirm the filled stripe phase in the ground state of the doped pure ($t^{\prime}=0$) Hubbard model. When the next nearest neighboring hopping term ($t^{\prime}=-0.2$) is included, the ground state is found to be half-filled stripes, consistent with experimental observation in cuprates. Noticeably, the direction of the stripe for $t^{\prime}=-0.2$ is determined to be horizontal when the width and the length are not equal, different from previous studies  (?). These results indicate previous calculations on systems with open boundary suffer from the boundary effect and underscoring the effectiveness of the new NQS method in gaining more reliable results.
We also have a careful analysis of the attention patterns learned in the ground state of 2D Hubbard model. Interestingly, we find that after the energy optimizations, different heads in the NQS ansatz specialize in encoding the correlations in different scales in the ground-state wave function. This special structure enables our NQS ansatz to effectively capture long-range correlations and entanglements, which is crutial for the accurate calculation of strongly correlated quantum many-body systems.
There is much room for the future development of NQS. On one hand, the development of more efficient optimization approach could enable the study of even large system sizes, which is crucial for more reliable finite size scaling in the doped Hubbard model. On the other hand, more sophisticated NQS ansatz for the calculation of dynamic and finite temperature properties  (?) are needed to further unveil the physics of the Hubbard model and establish the microscopic theory for cuprate. We anticipate that NQS framework will continue benefiting from the development in the broad machine learning community. Nevertheless, our results demonstrate that NQS is already a powerful framework for solving challenging many-fermions systems.
@@
In VMC, we employ Markov chain Monte Carlo (MCMC) to sample electron configurations according to $\psi_{\theta}^{2}(\mathbf{n})$. Following convergence, a separate inference run is performed with the fixed, optimized NQS to compute the physical observables until desired statistical accuracy is achieved. Further details on the network architecture, hyperparameters, and algorithmic specifics are provided in the supplementary materials.
@@ 1 Variational Monte Carlo
is the local energy. Exactly calculation of the energy using Equation S1 is infeasible in an exponentially large Hilbert space. However, we can approximate this summation using the Monte Carlo method. A set of occupation numbers $\{\mathbf{n}_{i}\}_{i=1}^{B}$ is sampled from the probability distribution defined by the squared amplitude of the wavefunction using Markov chain Monte Carlo (MCMC) techniques. The local energy of occupation numbers, $E_{\text{loc}}^{\theta}(\mathbf{n})$, can be efficiently calculated from the Hubbard model Hamiltonian, with the number of summation terms in Equation S2 scaling linearly with the size of the system. Therefore, parameters of the neural network can be optimized to approch ground state of Hubbard model:
@@ 1.1 Neural Quantum State Wavefunction Ansatz
$\text{Attn}^{(l)}(X)$ $=\sum_{h=1}^{H}\text{softmax}\left(XW_{Q}^{(l,h)}(XW_{K}^{(l,h)})^{\top}/\sqrt{d/H}\right)XW_{V}^{(l,h)}(W_{O}^{(l,h)})^{\top},$ (S6)
@@
The variational parameters $W_{Q}^{(l,h)},W_{K}^{(l,h)},W_{V}^{(l,h)},W_{O}^{(l,h)}\in\mathbb{R}^{d\times d/H}$ are the query, key, value, and output matrices of the head $h$, respectively. $W_{F}^{(l)}\in\mathbb{R}^{d\times d}$ is the weight matrix in the FFN(l). The activation $\sigma$ is chosen as SiLU  (?).
The final output $X^{(L)}$ is transformered by a spin-dependent linear layer to produce $K$ sets of backflow orbitals $M^{\sigma,k}\in\mathbb{R}^{N\times N_{e}}$, where $\sigma\in\{\uparrow,\downarrow\}$ and $k\in[K]$. The Slater matrix $\Phi^{k}$ is formed by selecting rows of $M^{\sigma,k}$ corresponding to the occupied position in the input configuration $n$. The overall NQS wavefunction is a linear combination of these determinants:
@@ 1.2 Optimization
Stochastic reconfiguration (SR) is one of the most popular optimization methods for NQS. In this work, however, we propose an advanced SR algorithms for superior performance, named Moment-Adaptive Reconfiguration Solver (MARCH), most directly inspired by the SPRING algorithm  (?). MARCH combines ideas from the extremely successful Adam optimizer in traditional machine learning  (?) and the recently introduced efficient Min-SR algorithm  (?).
In SR  (?, ?, ?), we optimize variational wavefunction |$\ket{\psi_{\theta}}$ by finding parameter updates that most approaching imaginary-time evolution wavefunction $e^{-\tau\hat{H}}\ket{\psi}$. As imaginary time evolution may push the state out of parameter manifold, SR utilizes Fubini-Study (FS) distance to measure the quantum distance and pull the imaginary-time evolving state back. In general, given two quantum states |$\ket{\psi},\ket{\phi}$, FS distance is defined as follows:
$d(\ket{\psi},\ket{\phi})=\text{arccos}\frac{\left|\braket{\psi}{\phi}\right|}{||\psi||\cdot||\phi||}$ (S9)
By expanding the FS distance to the lowest order, one derive:
$d(\ket{\psi}+\ket{\delta\psi},\ket{\phi}+\ket{\delta\phi})=\left|\left|\left(\frac{\ket{\delta\psi}}{||\psi||}-\frac{\ket{\delta\phi}}{||\phi||}\right)-\left(\frac{\braket{\psi}{\delta\psi}\ket{\psi}}{||\psi||^{3}}-\frac{\braket{\phi}{\delta\phi}\ket{\phi}}{||\phi||^{3}}\right)\right|\right|$ | | (S10)
@@
|$\ket{\delta\psi_{\theta}}=\sum\limits_{\mathbf{n}}\sum\limits_{k}\frac{\partial\psi_{\theta}(\mathbf{n})}{\partial\theta_{k}}\delta\theta_{k}\ket{\mathbf{n}}=\sum\limits_{\mathbf{n}}\psi_{\theta}(\mathbf{n})\sum\limits_{k}\bar{O}_{k}(\mathbf{n})\delta\theta_{k}\ket{\mathbf{n}}$ (S11)
|$\ket{\delta\psi^{\prime}}=-\hat{H}\delta\tau\ket{\psi_{\theta}}=-\delta\tau\sum\limits_{\mathbf{n}}\psi_{\theta}(\mathbf{n})E_{\text{loc}}^{\theta}(\mathbf{n})\ket{\mathbf{n}}$ (S12)
@@
$d^{2}(\ket{\psi_{\theta}}+\ket{\delta\psi_{\theta}},\ket{\psi_{\theta}}+\ket{\delta\psi^{\prime}})=\sum\limits_{\mathbf{n}}\frac{\psi^{2}_{\theta}(\mathbf{n})}{||\psi_{\theta}(\mathbf{n})||^{2}}\left|\sum\limits_{k}O_{k}(\mathbf{n})\delta\theta_{k}-\epsilon(\mathbf{n})\right|^{2}=\mathbb{E}_{\mathbf{n}}\left[\left|O(\mathbf{n})\cdot\delta\bm{\theta}-\epsilon(\mathbf{n})\right|^{2}\right]$ (S13)
where $O_{k}(\mathbf{n})=\bar{O}_{k}(\mathbf{n})-\mathbb{E}_{\mathbf{n}}[\bar{O}_{k}(\mathbf{n})],O(\mathbf{n})=\left(O_{1}(\mathbf{n}),\cdots,O_{k}(\mathbf{n}),\cdots\right)^{T},\epsilon(\mathbf{n})=-\delta\tau(E_{\text{loc}}(\mathbf{n})-\mathbb{E}_{\mathbf{n}}[E_{\text{loc}}(\mathbf{n})])$.
@@
$\mathrm{d}\theta=\operatorname*{arg\,min}_{\mathrm{d}\theta^{\prime}}\mathbb{E}_{\mathbf{n}}\left[|O(\mathbf{n})\mathrm{d}\theta^{\prime}-\epsilon(\mathbf{n})|^{2}\right]\approx\operatorname*{arg\,min}_{\mathrm{d}\theta^{\prime}}\frac{1}{B}\sum\limits_{i=1}^{B}|O(\mathbf{n}_{i})\mathrm{d}\theta^{\prime}-\epsilon(\mathbf{n}_{i})|^{2}$ (S14)
@@
To express the problem in a standard linear algebra form, we construct a matrix $\mathbf{O}$ and a vector $\bm{\epsilon}$ by stacking the contributions from each of the $B$ samples in the batch:
$\mathbf{O}=\begin{pmatrix}O(\mathbf{n}_{1})\\ O(\mathbf{n}_{2})\\ \vdots\\ O(\mathbf{n}_{B})\end{pmatrix}\in\mathbb{R}^{B\times N_{\text{params}}}\quad\text{and}\quad\bm{\epsilon}=\begin{pmatrix}\epsilon(\mathbf{n}_{1})\\ \epsilon(\mathbf{n}_{2})\\ \vdots\\ \epsilon(\mathbf{n}_{B})\end{pmatrix}\in\mathbb{R}^{B}$ (S15)
Since $\mathbf{O}$ might be ill-conditioned, the conventional method is to add some form of regularization. For example, a Tikhonov regularization is usually added to yield the regularized problem as follows:
$\mathrm{d}\theta=\operatorname*{arg\,min}_{\mathrm{d}\theta^{\prime}}\frac{1}{\lambda}||\mathbf{O}\mathrm{d}\theta^{\prime}-\bm{\epsilon}||^{2}+||\mathrm{d}\theta^{\prime}||^{2}$ (S16)
MARCH differs from previous SR algorithms by adaptively translates and scales $\mathrm{d}\theta$ individually for each parameter based on the estimation of the first and second moments.
@@
$\mathrm{d}\theta=\operatorname*{arg\,min}_{\mathrm{d}\theta^{\prime}}\frac{1}{\lambda}||\mathbf{O}\mathrm{d}\theta^{\prime}-\bm{\epsilon}||^{2}+||\text{diag}(V_{k-1})^{1/4}(\mathrm{d}\theta^{\prime}-\phi_{k-1})||^{2}$ (S17)
where $\text{diag}(V_{k-1})$ is the diagonal matrix whose main diagonal is composed of $V_{k-1}$.
The parameter updates are guided by two estimators. The first moment estimator, $\phi_{k}=\mu\mathrm{d}\theta_{k}$, acts as a momentum term. Intuitively, $\phi$ carries the momentum, which accelerates updates in directions of persistent gradients and dampens oscillations. The second moment estimator, $V_{k}=\beta V_{k-1}+(\mathrm{d}\theta_{k}-\mathrm{d}\theta_{k-1})^{2}$ , tracks the volatility of the gradient. A large $V_{k}$ indicates that the gradient is changing erratically, so the update of parameter is suppressed to prevent divergence. Conversely, a small $V_{k}$ implies a stable gradient, allowing the parameter to be updated more aggressively. The behavior of these estimators is governed by the hyperparameters $\mu$, $\beta$, and $\lambda$, which are specified in Table S3.
Therefore, the explicit update formula can be derived. We first define $\pi=\text{diag}(V_{k-1})^{1/4}(\mathrm{d}\theta^{\prime}-\phi_{k-1}),U=\mathbf{O}\text{diag}(V_{k-1})^{-1/4},\zeta=\bm{\epsilon}-\mathbf{O}\phi_{k-1}$. The equation S17 can be recast as:
@@
Via Woodbury matrix identity, the solution is given by:
@@
$\mathrm{d}\theta=\text{diag}(V_{k-1})^{-1/2}\mathbf{O}^{T}(\mathbf{O}\text{diag}(V_{k-1})^{-1/2}\mathbf{O}^{T}+\lambda I)^{-1}(\bm{\epsilon}-\mathbf{O}\phi_{k-1})+\phi_{k-1}$ (S23)

### v2

@@ Abstract
Figure 5: Energy difference between vertical stripe and horizontal stripe. Kinetic energy, potential energy, and total energy results are all shown. We systematically study systems $L_{x}\times 8$ with $L_{x}$ ranging from 8 to 32, with $t^{\prime}=-0.2$ and $\delta=1/8$. We can find that the horizontal stripe has lower energy for all the systems (notice that for $8\times 8$ system, vertical and horizontal stripes are the same). Moreover, the results show that the gain of energy in the horizontal stripe is mainly from the kinetic energy.
Previous numerical studies on stripe order in the Hubbard and related models have been conducted on cylindrical geometries with DMRG  (?, ?, ?, ?, ?). However, such boundary conditions and their related finite-size effect can inherently favor the formation of vertical stripe orientations. To alleviate the possible boundary effect, we study systems with PBC in both directions. We systematically study systems with size $L_{x}\times 8$ where $L_{x}$ is ranged from $8$ to $32$, with $t^{\prime}=-0.2$ and $\delta=1/8$ (see Equation 1). We consistently observe that a horizontal stripe configuration is energetically favored in the ground state while the system can accommodate half-filled stripe in both horizontal and vertical directions (see Figure 4(B) and (C), Figure 5 and S21-S26 in the Supplementary Materials). These results suggest the existence of a tendency toward fewer long stripes rather than multiple fragmented ones in the systems, different from previous studies on cylinders  (?). From Figure 5, we can also find that the gain of energy for the horizontal stripe state is mainly from the kinetic energy, while the potential energy in the horizontal stripe is slightly higher. A simple explanation is that the formation of a long “river” of holes in the horizontal stripe is beneficial to the kinetic energy. A comprehensive understanding of this observation needs further investigation.
For rectangular systems with a width not divisible by $8$, a horizontal stripe with wavelength $\lambda=4$ is impossible, so “frustration” appears. One such example is the $16\times 10$ system, for which we find that the trend for horizontal stripe wins and the ground state has a wavelength of $\lambda=5$ while the vertical stripe state with $\lambda=4$ has a higher energy (see Figure S28-S29 in the Supplementary Materials). In another example for the $16\times 12$ system, the trend for wavelength $4$ wins in the competition and the ground state is a vertical stripe (see Figure S30 in the Supplementary Materials).
For the pure Hubbard model, it is also possible that the true ground state is a horizontal stripe. But the resolution of this issue requires the study of systems that can accommodate stripes with a wavelength $8$ in both vertical and horizontal directions. The smallest size for such systems under PBC is $32\times 16$, which is beyond our present capacity.
@@
In this work, we propose a new NQS framework for studying quantum many-body systems by leveraging the cutting-edge machine learning architectures and developing highly accurate optimization algorithms. State-of-the-art results for the 2D doped Hubbard model with unprecedented sizes are achieved with this new framework. NQS enables the study of 2D Hubbard model under periodic boundary conditions, which is crucial to get rid of the boundary effect and makes the extrapolation to the thermodynamic limit reliable. By studying 2D Hubbard model with large systems under periodic boundary conditions, we confirm the filled stripe phase in the ground state of the doped pure ($t^{\prime}=0$) Hubbard model. When the next nearest neighboring hopping term ($t^{\prime}=-0.2$) is included, the ground state is found to be half-filled stripes, consistent with experimental observation in cuprates. Noticeably, the direction of the stripe for $t^{\prime}=-0.2$ is determined to be horizontal when the width and the length are not equal, different from previous studies  (?). These results indicate that previous calculations on systems with open boundary suffer from the boundary effect and underscore the effectiveness of the new NQS method in gaining more reliable results.
We also have a careful analysis of the attention patterns learned in the ground state of 2D Hubbard model. Interestingly, we find that after the energy optimizations, different heads within our NQS specialize in encoding the correlations at different scales of the ground-state wave function. This special structure enables our NQS ansatz to effectively capture long-range correlations and entanglements, which is crucial for the accurate calculation of strongly correlated quantum many-body systems.
There is much room for the future development of NQS. On one hand, the development of more efficient optimization approaches could enable the study of even large system sizes, which is crucial for more reliable finite size scaling in the doped Hubbard model. On the other hand, more sophisticated NQS ansatz for the calculation of dynamic and finite temperature properties  (?) are needed to further unveil the physics of the Hubbard model and establish the microscopic theory for cuprate. We anticipate that NQS framework will continue benefiting from the development in the broad machine learning community. Nevertheless, our results demonstrate that NQS is already a powerful framework for solving challenging many-fermions systems.
@@
In VMC, we employ Markov chain Monte Carlo (MCMC) to sample electron configurations according to $\psi_{\theta}^{2}(\mathbf{n})$. Following convergence, a separate inference run is performed with the fixed, optimized NQS to compute the physical observables until the desired statistical accuracy is achieved. Further details on the network architecture, hyperparameters, and algorithmic specifics are provided in the supplementary materials.
@@ Acknowledgements
Data and materials availability:
All code and data will be made openly available upon publication.
@@
Yuntian Gu, Wenrui Li, Heng Lin, Bo Zhan, Ruichen Li, Yifei Huang,
Di He, Yantao Wu, Tao Xiang, Mingpu Qin, Liwei Wang, Dingshun Lv
Corresponding author: dihe@pku.edu.cn (D. He); yantaow@iphy.ac.cn (Y. T. Wu);
txiang@iphy.ac.cn (T. Xiang); qinmingpu@sjtu.edu.cn(M. P. Qin);
wanglw@pku.edu.cn (L. W. Wang); lvdingshun@bytedance.com (D. S. Lv).
These authors contributed equally to this work.
@@ 1 Variational Monte Carlo
is the local energy. Exact calculation of the energy using Equation S1 is infeasible in an exponentially large Hilbert space. However, we can approximate this summation using the Monte Carlo method. A set of occupation numbers $\{\mathbf{n}_{i}\}_{i=1}^{B}$ is sampled from the probability distribution defined by the squared amplitude of the wavefunction using Markov chain Monte Carlo (MCMC) techniques. The local energy of occupation numbers, $E_{\text{loc}}^{\theta}(\mathbf{n})$, can be efficiently calculated from the Hubbard model Hamiltonian, with the number of summation terms in Equation S2 scaling linearly with the size of the system. Therefore, parameters of the neural network can be optimized to approachthe ground state of the Hubbard model:
@@ 1.1 Neural Quantum State Wavefunction Ansatz
$\text{Attn}^{(l)}(X)$ $=\sum_{h=1}^{H}\text{softmax}\left(XW_{Q}^{(l,h)}(XW_{K}^{(l,h)})^{\top}/\sqrt{d_{H}}\right)XW_{V}^{(l,h)}(W_{O}^{(l,h)})^{\top},$ (S6)
@@
The variational parameters $W_{Q}^{(l,h)},W_{K}^{(l,h)},W_{V}^{(l,h)},W_{O}^{(l,h)}\in\mathbb{R}^{d\times d_{H}}$ are the query, key, value, and output matrices of the head $h$, respectively. $d_{H}=d/H$ is the hidden dimension of each head. $W_{F}^{(l)}\in\mathbb{R}^{d\times d}$ is the weight matrix in the FFN(l). The activation $\sigma$ is chosen as SiLU  (?).
The final output $X^{(L)}$ is transformered by a linear layer to produce $K$ sets of backflow orbitals $M^{k}\in\mathbb{R}^{2N\times N_{e}}$, where $k\in[K]$. The Slater matrix $\Phi^{k}$ is formed by selecting rows of $M^{k}$ corresponding to the occupied position in the input configuration $\mathbf{n}$. The overall NQS wavefunction is a linear combination of these determinants:
@@ 1.2 Optimization
Stochastic reconfiguration (SR) is one of the most popular optimization methods for NQS. In this work, however, we propose an advanced SR algorithms for superior performance, named Moment-Adaptive ReConfiguration Heuristic (MARCH), most directly inspired by the SPRING algorithm  (?). MARCH combines ideas from the extremely successful Adam optimizer in traditional machine learning  (?) and the recently introduced efficient Min-SR algorithm  (?, ?).
In SR, we optimize variational wavefunction |$\ket{\psi_{\theta}}$ by finding parameter updates that best approach imaginary-time evolution wavefunction $e^{-\tau\hat{H}}\ket{\psi}$. As imaginary time evolution may push the state out of the parameter manifold, SR utilizes the Fubini-Study (FS) distance to measure the quantum distance and pull the imaginary-time evolving state back. In general, given two quantum states |$\ket{\psi},\ket{\phi}$, FS distance is defined as follows:
$d(\ket{\psi},\ket{\phi})=\text{arccos}\frac{\left|\braket{\psi}{\phi}\right|}{||\psi||\cdot||\phi||}.$ (S9)
By expanding the FS distance to the lowest order, one derives:
$d(\ket{\psi}+\ket{\delta\psi},\ket{\psi}+\ket{\delta\phi})=\left|\left|\left(\frac{\ket{\delta\psi}}{||\psi||}-\frac{\ket{\delta\phi}}{||\psi||}\right)-\left(\frac{\braket{\psi}{\delta\psi}\ket{\psi}}{||\psi||^{3}}-\frac{\braket{\psi}{\delta\phi}\ket{\psi}}{||\psi||^{3}}\right)\right|\right|.$ | | . (S10)
@@
|$\ket{\delta\psi_{\theta}}=\sum\limits_{\mathbf{n}}\sum\limits_{k}\frac{\partial\psi_{\theta}(\mathbf{n})}{\partial\theta_{k}}\delta\theta_{k}\ket{\mathbf{n}}=\sum\limits_{\mathbf{n}}\psi_{\theta}(\mathbf{n})\sum\limits_{k}\bar{O}_{k}(\mathbf{n})\delta\theta_{k}\ket{\mathbf{n}},$ (S11)
|$\ket{\delta\psi^{\prime}}=-\hat{H}\delta\tau\ket{\psi_{\theta}}=-\delta\tau\sum\limits_{\mathbf{n}}\psi_{\theta}(\mathbf{n})E_{\text{loc}}^{\theta}(\mathbf{n})\ket{\mathbf{n}},$ (S12)
@@
$d^{2}(\ket{\psi_{\theta}}+\ket{\delta\psi_{\theta}},\ket{\psi_{\theta}}+\ket{\delta\psi^{\prime}})=\sum\limits_{\mathbf{n}}\frac{\psi^{2}_{\theta}(\mathbf{n})}{||\psi_{\theta}||^{2}}\left|\sum\limits_{k}O_{k}(\mathbf{n})\cdot\delta\theta_{k}-\epsilon(\mathbf{n})\right|^{2},$ (S13)
where $O_{k}(\mathbf{n})=\bar{O}_{k}(\mathbf{n})-\mathbb{E}_{\mathbf{n}^{\prime}}[\bar{O}_{k}(\mathbf{n}^{\prime})],O(\mathbf{n})=\left(O_{1}(\mathbf{n}),\cdots,O_{k}(\mathbf{n}),\cdots\right)^{T},\epsilon(\mathbf{n})=-\delta\tau(E_{\text{loc}}(\mathbf{n})-\mathbb{E}_{\mathbf{n}^{\prime}}[E_{\text{loc}}(\mathbf{n}^{\prime})])$.
@@
$\mathrm{d}\theta=\operatorname*{arg\,min}_{\mathrm{d}\theta^{\prime}}\mathbb{E}_{\mathbf{n}}\left[|O(\mathbf{n})\mathrm{d}\theta^{\prime}-\epsilon(\mathbf{n})|^{2}\right]\approx\operatorname*{arg\,min}_{\mathrm{d}\theta^{\prime}}\frac{1}{B}\sum\limits_{i=1}^{B}|O(\mathbf{n}_{i})\mathrm{d}\theta^{\prime}-\epsilon(\mathbf{n}_{i})|^{2},$ (S14)
@@
To express the problem in a standard linear algebra form, we construct a matrix $\tilde{O}$ and a vector $\tilde{\epsilon}$ by stacking the contributions from each of the $B$ samples in the batch:
$\tilde{O}=\begin{pmatrix}O(\mathbf{n}_{1})\\ O(\mathbf{n}_{2})\\ \vdots\\ O(\mathbf{n}_{B})\end{pmatrix}\in\mathbb{R}^{B\times N_{\text{params}}}\quad\text{and}\quad\tilde{\epsilon}=\begin{pmatrix}\epsilon(\mathbf{n}_{1})\\ \epsilon(\mathbf{n}_{2})\\ \vdots\\ \epsilon(\mathbf{n}_{B})\end{pmatrix}\in\mathbb{R}^{B}.$ (S15)
Since $\tilde{O}$ might be ill-conditioned, the conventional method is to add some form of regularization. For example, a Tikhonov regularization is usually added to yield the regularized problem as follows:
$\mathrm{d}\theta=\operatorname*{arg\,min}_{\mathrm{d}\theta^{\prime}}\frac{1}{\lambda}||\tilde{O}\mathrm{d}\theta^{\prime}-\tilde{\epsilon}||^{2}+||\mathrm{d}\theta^{\prime}||^{2}.$ (S16)
MARCH differs from previous SR algorithms by adaptively translating and scaling $\mathrm{d}\theta$ individually for each parameter based on the estimation of the first and second moments.
@@
$\mathrm{d}\theta_{k}=\operatorname*{arg\,min}_{\mathrm{d}\theta^{\prime}}\frac{1}{\lambda}||\tilde{O}\mathrm{d}\theta^{\prime}-\tilde{\epsilon}||^{2}+||\text{diag}(v_{k-1})^{1/4}(\mathrm{d}\theta^{\prime}-\phi_{k-1})||^{2},$ (S17)
where $k$ denotes the current step, $\text{diag}(v_{k-1})$ is the diagonal matrix whose main diagonal is composed of $v_{k-1}$. The first-order moment estimator, $\phi_{k}=\mu\mathrm{d}\theta_{k}$, acts as a momentum term. Intuitively, $\phi$ carries the momentum, which accelerates updates in directions of persistent gradients and dampens oscillations. The second moment estimator, $v_{k}=\beta v_{k-1}+(\mathrm{d}\theta_{k}-\mathrm{d}\theta_{k-1})^{2}$ , tracks the volatility of the gradient. A large $v_{k}$ indicates that the gradient is changing erratically, so the update of the parameter is suppressed to prevent divergence. Conversely, a small $v_{k}$ implies a stable gradient, allowing the parameter to be updated more aggressively. The behavior of these estimators is governed by the hyperparameters $\mu$, $\beta$, and $\lambda$, which are specified in Table S7.
Therefore, the explicit update formula can be derived. We first define $\pi=\text{diag}(v_{k-1})^{1/4}(\mathrm{d}\theta^{\prime}-\phi_{k-1}),U=\tilde{O}\text{diag}(v_{k-1})^{-1/4},\zeta=\tilde{\epsilon}-\tilde{O}\phi_{k-1}$. The equation S17 can be recast as:
@@
We can reduce the computational budget from $\mathcal{O}(N_{\text{params}}^{3})$ to $\mathcal{O}(B^{2}N_{\text{params}})$ by Woodbury matrix identity:
@@
$\mathrm{d}\theta=\text{diag}(v_{k-1})^{-1/2}\tilde{O}^{T}(\tilde{O}\text{diag}(v_{k-1})^{-1/2}\tilde{O}^{T}+\lambda I)^{-1}(\tilde{\epsilon}-\tilde{O}\phi_{k-1})+\phi_{k-1}.$ (S23)